# Diffusion Models for Robotic Manipulation: A Survey

### v2

@@ 5 Experiments and Benchmarks
In this section, we focus on the evaluation of the various DMs for robotic manipulation. Details on the employed benchmarks and baselines are listed in separate tables for imitation learning (Table 6), reinforcement learning (Table 7), and grasp learning (Table 5). Separately, the references for all applied benchmarks are listed in Table 9.
@@
Reuss et al., (2024) CALVIN, LIBERO 24h, 50 4.5h
@@
Chen et al., (2024) DP, SE(3)-DM FrankaKitchen, Adroit 16k - 64k, 1.25k - 5k 60
@@
Many methods are only being evaluated against baselines, which are not based on DMs themselves. However, there are some common DM-based baselines. For methods operating in $\mathbf{SE}(3)$-space (Chen et al.,, 2024; Song et al., 2024b, ; Ryu et al.,, 2024), $\mathbf{SE}(3)$-Diffusion Policy (Urain et al.,, 2023), probably the first paper using DMs for grasp generation, is commonly used as baseline. For RL-based methods, the RL-based Diffuser (Janner et al.,, 2022), Diffusion-QL (Wang et al., 2023b, ), and Decision Diffuser (Ajay et al.,, 2023) are commonly used as baselines. It should be noted that in the original paper, Decision Diffuser (Ajay et al.,, 2023) is evaluated against Diffuser (Janner et al.,, 2022) and outperforms it on almost all tasks, particularly on the manipulation tasks, block stacking, and rearrangement. However, neither of these methods is evaluated on real-world tasks. Another common baseline is DP (Chi et al.,, 2023), as many methods are developed based on it. A common baseline for methods integrating 3D visual representations is 3D Diffusion Policy(Ze et al.,, 2024). 3D Diffusion Policy is evaluated against DP, and outperforms it on a huge variety of tasks in the benchmarks Adroit, MetaWorld, and Dexart with an average success rate of 74.4%, outperforming DP by 24.2%. It is also evaluated on four real-world manipulation tasks: rolling and pinching a dumpling, drilling, and pouring. With an average success rate of 85.0% it outperforms DP by 50%. 3D Diffusion Policy is greatly outperformed by 3D Diffuser Actor (Ke et al.,, 2024) on the CALVIN benchmark, especially for zero-shot long-horizon tasks. However, no comparison for real-world tasks is provided.
6 Conclusion, Limitations and Outlook
6.1 Limitations
6.1.1 Generalizability
6.1.2 Sampling speed
The principal limitation inherent to DMs can be attributed to the iterative nature of the sampling process, which results in a time-intensive sampling procedure, thus impeding efficiency and real-time prediction capabilities. Despite recent advances that improve sampling speed and quality (Chen et al.,, 2024; Zhou et al., 2024a, ), a considerable number of recent methods use DDIM (Song et al., 2021a, ), although other methods, such as DPM-solver (Lu et al.,, 2022) have shown better performance. However, this comparison has only been performed using image generation benchmarks and would need to be verified for applications in robotic manipulation. Numerous works have demonstrated competitive task performance using DDIM, but do not directly investigate the decrease in task performance associated with a lower number of reverse diffusion steps. Ko et al., (2024) analyzes their approach using both DDPM and DDIM sampling, reporting a sampling process that is ten times faster with only a 5.6% decrease in task performance when using DDIM. Although such a decline might appear negligible, its significance is highly task-dependent. Consequently, there is a need for efficient sampling strategies and a more comprehensive analysis of existing sampling methods, particularly regarding the domain of robotic manipulation. It should, however, be noted that already in DP (Chi et al.,, 2023), one of the earlier methods combining DMs with receding-horizon control for trajectory planning, real-time control is possible. Using DDIM with 10 denoising steps during inference, they report an inference latency of 0.1s on a Nvidia 3080 GPU.
6.2 Conclusion and Outlook
A key challenge of DMs is the slow inference speed. In the field of computer vision, fast samplers have been developed that have not yet been evaluated in the field of robotic manipulation. Testing those samplers and comparing them against the commonly used ones, could be one step to increase sampling efficiency. Moreover, there are also methods for fast sampling, specifically in robotic manipulation, that are not broadly used, e.g. BRIDGeR (Chen et al.,, 2024). While the generalizability of DMs remains also an open challenge, the image generation capabilities of DMs open new avenues in data augmentation for data scaling, making methods more robust to limited data variety. Generalizability could be also improved by the integration of advanced vision-language, and vision-language action models.
Acknowledgment
We thank our colleague Edgar Welte for providing the video data of for the illustration of the diffusion process in Fig. 2.
Funding
Appendix

### v3

@@ 5 Experiments and Benchmarks
In this section, we focus on the evaluation of the various DMs for robotic manipulation. Details on the employed benchmarks and baselines are listed in the separate tables for imitation learning (Table 6), reinforcement learning (Table 7) in the Appendix, and grasp learning (Table 5). Separately, the references for all applied benchmarks are listed in Table 9 in theAppendix.
Many methods are only being evaluated against baselines, which are not based on DMs themselves. However, there are some common DM-based baselines. For methods operating in $\mathbf{SE}(3)$-space (Chen et al., 2024a, ; Song et al., 2024b, ; Ryu et al.,, 2024), $\mathbf{SE}(3)$-Diffusion Policy (Urain et al.,, 2023), probably the first paper using DMs for grasp generation, is commonly used as baseline. For RL-based methods, the RL-based Diffuser (Janner et al.,, 2022), Diffusion-QL (Wang et al., 2023b, ), and Decision Diffuser (Ajay et al.,, 2023) are commonly used as baselines. It should be noted that in the original paper, Decision Diffuser (Ajay et al.,, 2023) is evaluated against Diffuser (Janner et al.,, 2022) and outperforms it on almost all tasks, particularly on the manipulation tasks, block stacking, and rearrangement. However, neither of these methods is evaluated on real-world tasks. Another common baseline is DP (Chi et al.,, 2023), as many methods are developed based on it. A common baseline for methods integrating 3D visual representations is 3D Diffusion Policy(Ze et al.,, 2024). 3D Diffusion Policy is evaluated against DP, and outperforms it on a huge variety of tasks in the benchmarks Adroit, MetaWorld, and Dexart with an average success rate of 74.4%, outperforming DP by 24.2%. It is also evaluated on four real-world manipulation tasks: rolling and pinching a dumpling, drilling, and pouring. With an average success rate of 85.0% it outperforms DP by 50%. 3D Diffusion Policy is greatly outperformed by 3D Diffuser Actor (Ke et al.,, 2024) on the CALVIN benchmark, especially for zero-shot long-horizon tasks. However, no comparison for real-world tasks is provided.
6 Conclusion, Limitations and Outlook
6.1 Limitations
6.1.1 Generalizability
6.1.2 Sampling speed
The principal limitation inherent to DMs can be attributed to the iterative nature of the sampling process, which results in a time-intensive sampling procedure, thus impeding efficiency and real-time prediction capabilities. Despite recent advances that improve sampling speed and quality (Chen et al., 2024a, ; Zhou et al., 2024a, ), a considerable number of recent methods use DDIM (Song et al., 2021a, ), although other methods, such as DPM-solver (Lu et al.,, 2022) have shown better performance. However, this comparison has only been performed using image generation benchmarks and would need to be verified for applications in robotic manipulation. Numerous works have demonstrated competitive task performance using DDIM, but do not directly investigate the decrease in task performance associated with a lower number of reverse diffusion steps. Ko et al., (2024) analyzes their approach using both DDPM and DDIM sampling, reporting a sampling process that is ten times faster with only a 5.6% decrease in task performance when using DDIM. Although such a decline might appear negligible, its significance is highly task-dependent. Consequently, there is a need for efficient sampling strategies and a more comprehensive analysis of existing sampling methods, particularly regarding the domain of robotic manipulation. It should, however, be noted that already in DP (Chi et al.,, 2023), one of the earlier methods combining DMs with receding-horizon control for trajectory planning, real-time control is possible. Using DDIM with 10 denoising steps during inference, they report an inference latency of 0.1s on a Nvidia 3080 GPU.
6.2 Conclusion and Outlook
A key challenge of DMs is the slow inference speed. In the field of computer vision, fast samplers have been developed that have not yet been evaluated in the field of robotic manipulation. Testing those samplers and comparing them against the commonly used ones, could be one step to increase sampling efficiency. Moreover, there are also methods for fast sampling, specifically in robotic manipulation, that are not broadly used, e.g. BRIDGeR (Chen et al., 2024a, ). While the generalizability of DMs remains also an open challenge, the image generation capabilities of DMs open new avenues in data augmentation for data scaling, making methods more robust to limited data variety. Generalizability could be also improved by the integration of advanced vision-language, and vision-language action models.
Acknowledgment
We thank our colleague Edgar Welte for providing the video data for the illustration of the diffusion process in Fig. 2.
Funding
Appendix
@@ Acknowledgment
Reuss et al., 2024a CALVIN, LIBERO 24h, 50 4.5h
@@
Chen et al., 2024a DP, SE(3)-DM FrankaKitchen, Adroit 16k - 64k, 1.25k - 5k 60
@@
Reuss et al., 2024b DP CALVIN, LIBERO, Relay Kitchen,
Block Push
22966, 50, 566,
1000